OpenStudy (anonymous):

solve the initial problem dy/dt=1/(ty+t+y+1), y(0)=4

OpenStudy (anonymous):

I'd probably say use separation of variables :-) and note that:\[ty+t+y+1=t(y+1)+(y+1)=(t+1)(y+1)\]